# Selecting samples of galaxies with fewer Fingers-of-God

### v2

@@ Abstract
The radial positions of galaxies inferred from their measured redshift appear distorted due to their peculiar velocities. We argue that the contribution from stochastic velocities — which gives rise to ‘Fingers-of-God’ (FoG) anisotropy in the inferred maps — does not lend itself to perturbative modelling already on scales targeted by current experiments. To get around this limitation, we propose to remove FoG using data-driven indicators of their abundance that are local in nature and thus avoid selection biases. In particular, we show that the scale where the measured power spectrum quadrupole changes sign is tightly anti-correlated with both the satellite fraction and the velocity dispersion, and can thus be used to select galaxy samples with fewer FoG. In addition, we show that maps of the thermal Sunyaev-Zel’dovich distortion of the cosmic microwave background frequency spectrum can be used to identify and discard many of the most problematic galaxies. These techniques could potentially extend the reach of perturbative models for galaxy clustering and improve reconstructions of the large-scale velocity and displacement fields from the redshift-space positions of galaxies.
@@ 1 Introduction
We also explore the possibility of removing galaxies living in regions with strong thermal Sunyaev-Zel’dovich [33] or X-ray emission, as this indicates that those galaxies are located in massive halos where the thermal velocities of galaxies are large. We showcase the potential of this approach and emphasize the powerful complementarity between upcoming galaxy redshift surveys and CMB experiments such as the Simons Observatory (SO) [34] and CMB-S4 [35].
@@ 4.1 Breakdown of contributions to $P_{\ell}(k)$
According to our discussion around figure 1, the contribution from central-satellite and satellite-satellite pairs potentially becomes non-perturbative when the quadrupole of their power spectrum goes negative. In order to determine exactly what scale this corresponds to in either case, and also to understand the relative importance of different contributions to the measured multipoles, we will use the Abacus mocks again, populated with galaxies using a fixed HOD model — concretely, the ‘baseline’ model of Ref. [46] with the best-fit value from analyzing BOSS CMASS data; the only change we introduce is to set the completeness to $f_{\mathrm{ic}}=1$ to avoid additional modeling difficulties [^7] All in all, the HOD parameters are the following: $\log_{10}(M_{1}/h^{-1}\,M_{\odot})=14.10$, $\kappa=0.2$, $\alpha=1.12$, $\log_{10}(M_{\rm cut}/h^{-1}\,M_{\odot})=12.86$, $\log_{10}\sigma=-2.8$, $\alpha_{\rm c}=0.22$, $\alpha_{\rm s}=0.98$ and $f_{\rm ic}=1$.. This time, however, we will use the weights in equation (3.5) to disentangle contributions from central-central, central-satellite and satellite-satellite pairs.
@@ 4.2.2 An analysis of Abacus mocks with varying satellite fractions
Since the satellite fraction correlates strongly with FoG strength, these results illustrate the difficulty of accounting for FoG in perturbative models such as the EFTofLSS. However, the reader will have noted that even in the case where we have no satellites, cosmological constraints become biased by $k_{\rm max}=0.25\,h\,\rm{Mpc}^{-1}$. This is in qualitative agreement with earlier work in Ref. [15] which showed, using $N$-body simulations populated with emission-line galaxies following an HOD, that lowering the satellite fraction leads to only slightly more accurate constraints on cosmological parameters with no significant gains in terms of the $k$-reach of the one-loop theory. This could well be because higher-loop contributions independent of FoG are becoming relevant. Unlike FoG, these corrections may well be perturbative; for example, Ref. [64] showed recently that two-loop corrections to the redshift-space matter power spectrum at $z=0.5$ are significant already at $k=0.15\,h\,\rm{Mpc}^{-1}$. In the next section, we set out to disentangle the impact of FoG from these other, higher-loop effects.
@@ 4.2.3 An analysis of a manifestly perturbative data vector acted on by FoG
We now use the central value of these best-fit parameters to draw new central-central, central-satellite and satellite spectra from the EFT model and scale them by FoG damping factors as required; finally we linearly combine the terms weighting each one of them by the appropriate factors of $f_{\rm sat}$. The resulting power spectrum multipoles for the full sample — extracted from the redshift-space power spectrum of equation (4.7) using (2.5) — are shown in figure 8. We also compare them to measurements on the Abacus mocks, seeing that our fitting ensures good qualitative agreement. Our synthetic data vector thus contains a mixture of central and satellite contributions that is representative of the truth (insofar as the simulations are realistic), but crucially has no two-loop contributions besides those introduced by FoG.
@@
The more relevant question is therefore whether cosmological constraints are biased when FoG are present. This subsection addresses this point in the limit that our synthetic data vector is made up exclusively of satellites (we will consider the ‘full sample’ in the next subsection). To do this, we will once again sample the posterior probability on the parameters given the synthetic data, using the priors in table 2, with only slight modifications to the likelihood described in §4.2.2. We once again use the Effort emulator for the velocileptors model, but since the synthetic data vector is drawn directly in terms of $k$ and $\mu$, there is no need to account for AP distortions (as they were never introduced into the data vector). Although the data vector is known with no uncertainty, it is still important to assign the correct relative weight between different $k$’s when fitting. We thus use an analytic disconnected covariance described in appendix B, assuming there is no shot noise in our measurements. The high accuracy of the Effort emulator and the fact that our data vector is noiseless allow us to rescale the covariance to an effective volume of $1000\,\rm{Gpc}^{3}$ while keeping emulator error limited to at most $0.5\,\sigma$ at each $k$. Working with such a large effective volume lets us avoid projection effects and dramatically reduces the uncertainty in our determination of the inference biases associated with FoG. Also, with such a large volume, the prior becomes relatively unimportant [67], and changes to it do not significantly impact our results.
@@
It is instructive to investigate the direction in which FoG drive the EFT parameters. In figure 10, we show the marginalized constraints on the nuisance parameters corresponding to the inferences above. In addition to the posterior mean and 68% credible intervals, we overlay also the prior. The reader may recall that the prior on the counterterm parameters was chosen so as to ensure that the EFT contributions remain a small fraction of the linear-theory contribution at $k_{\rm max}$; more specifically, one prior standard deviation away from zero, $\alpha_{n}$ constitutes a 50% correction to the linear-theory prediction for the $n$-th moment [22]. From figure 10, we learn that when FoG are present, the preferred counterterm parameter values lie in regions that are disfavored by the prior. $\alpha_{4}$, in particular, lies more than one standard deviation away from the prior center already at $k_{\rm max}=0.1\,h\rm{Mpc}^{-1}$ whether or not the hexadecapole is included. This points to a breakdown of perturbativity roughly at the scale where the quadrupole crosses zero. Note, however, that all counterterms respond: this is because stochastic velocities introduce a number of counterterms into the theory, all but one of which are fully degenerate with the redshift-space counterterms.
@@
At this point, it is worth noting that the lack of AP distortions and two-loop effects in this toy model makes a direct comparison between figures 11 and 7 unwarranted. The lack of AP, in particular, means that our exercise is not sensitive to the ways in which the anisotropy induced by FoG can interfere with the extraction of cosmological information from the AP effect; this is expected to impact the $\omega_{c}$ and $H_{0}$ parameters primarily. Nevertherless, we do note that the trend seen in figure 7 where going from $f_{\rm sat}=0$ to $f_{\rm sat}=0.11$ shifts $\Delta\omega_{c}$ and $\Delta H_{0}$ in the negative direction and $\Delta A_{s}$ in the positive is reproduced also in figure 11 as the satellite fraction is increased.
Once again, it is worth visualizing the constraints on nuisance parameters corresponding to these inferences. These are shown in figure 12. The striking growth in $b_{1}$ with the satellite fraction is due to the mock sample having a higher proportion of satellites, which have higher bias. More interestingly, the $\alpha_{n}$ counterterms absorb FoG contributions and become large compared the standard deviation of the prior imposed on them, which suggests that the counterterms are no longer a small correction to the linear theory contribution. This is important for two reasons: first, it tells us that the perturbative expansion is breaking down; and second, in current data analyses, the priors imposed on the counterterms are typically chosen to be informative in order to avoid prior volume effects. For a given prior on $\alpha_{n}$, mischaracterizing the impact of FoG contributions to this parameter can effectively shrink the prior on contributions from other processes and potentially bias cosmological constraints.
@@ 5.1 Discarding the reddest galaxies
The connection we showed between the zero-crossing of the quadrupole and the prevalence of FoG suggests using this observable as a data-driven and model-independent way to select samples less affected by FoG. This connection draws attention to the findings of Ref. [68] (hereafter SYS21), who showed that the zero-crossing of the quadrupole is very sensitive to the color of galaxies (see their figure 15), and, to a lesser extent, also to their luminosity (their figure 14). Concretely, they showed that the power spectrum quadrupole of the reddest 20% of BOSS LOWZ galaxies crosses zero at $k\approx 0.2\,h\,\rm{Mpc}^{-1}$, while the full sample has a crossing at $k\approx 0.3\,h\,\rm{Mpc}^{-1}$. This suggests that splitting samples in color and simply dropping the reddest quintile could help mitigate FoG. Since the color selection is ‘local’ (in the sense that it is sensitive to the properties of individual galaxies rather than the large-scale structure of the Universe), this selection should not introduce the kinds of modeling complications that plague alternative approaches to FoG removal.
Guided by the zero-crossing of the quadrupole, we have in fact ‘rediscovered’ a relationship that is well-known in the literature: red galaxies display stronger FoG [29, 30, 31, 32]. This is related to the fact that stochastic velocities originate primarily from galaxies living in the most massive halos, and such high density regions are predominantly populated by redder galaxies, whereas bluer ones tend to avoid such environments [69, 70].
Let us now investigate how the zero-crossing of the quadrupole responds to discarding the reddest galaxies to produce a ‘bluer’ subsample. Since the Abacus simulations do not mock up the color of galaxies, we set out to approximate the different subsamples more roughly by trying to reproduce the power spectra reported in SYS21.
@@
Figure 13: Power spectrum multipoles illustrating the effect of removing the reddest 20% of galaxies from a BOSS-LOWZ-like sample. (Unlike in other plots in this work, we do not subtract shot noise from the monopole.) By reweighting galaxies to have an effective $f_{\rm sat}=0.3$ ($f_{\rm sat}=0.15$), we mock up a sample with a zero-crossing of the quadrupole at $k=0.21$ ($k=0.31$) resembling that of the reddest quintile (full sample) of BOSS LOWZ [68]. From this we can predict the spectra of the remaining 80% of galaxies after the reddest quintile is discarded, which has a zero-crossing at a higher $k=0.36$. This suggests that the remaining sample has a significantly smaller scale of FoG non-linearity.
@@
Figure 14: Marginalized posterior mean and 68% credible intervals on cosmological parameters as a function of $k_{\rm max}$ before (purple) and after (yellow) discarding the reddest quintile of mock galaxies resembling the BOSS LOWZ selection.
@@ 5.2 Masking galaxy clusters and groups
In §3, we showed that FoG are sourced primarily by galaxies living in the most massive halos, where deep gravitational potential wells source large virial velocities. These are the same clusters that, in turn, produce the strongest X-ray emission and Sunyaev-Zel’dovich (tSZ) distortion of the cosmic microwave background frequency spectrum. tSZ and X-ray intensity are tightly correlated with halo mass, so masking on these observables — for example, in regions with SNR of 5 or higher — could potentially remove the galaxies with the strongest peculiar velocities, and do so with a very clean selection function that depends only on halo mass and crucially not on galaxy overdensity or velocity. This is due in large part to the fact that structures that can be detected on an individual basis in current or upcoming Compton $y$-maps are all in the one-halo regime [^14] The impact of two-halo contributions to the $y$-maps can be addressed in hydro simulations, though we defer this to future work. (two-halo contributions to the tSZ effect are only evident after stacking very many objects). Moreover, these objects are typically identified only after high-pass filtering the Compton $y$-maps to eliminate any long-wavelength modes tracing the cosmic web. This means that the key discerning property in this selection is halo mass. The resulting galaxy sample will simply have a different (smaller) galaxy bias than the initial population; bias can of course be marginalized over in the standard ways, at least in the limit applicable here where only a relatively small fraction of the galaxies are masked [^15] There is likely a connection between what we propose to do and density-split statistics or marked power spectra, though we claim that our selection is more local in nature as it is guided by halo mass. Our approach will thus produce populations with clustering properties that are easier to model. The modeling will also be made easier by the fact that we typically identify — and discard — a very small number of objects.. Below, we will in fact verify that thresholding on halo mass does not introduce any noticeable biases on parameter inference; much on the contrary, it helps recover more accurate cosmological parameters.
As an example, suppose we are to analyze a galaxy redshift survey that fully overlaps with a complementary CMB survey. The tSZ signal being independent of redshift, we could in principle have a complete census of all the clusters above a certain mass threshold, and discard galaxies living within them [^16] Future work will be required to understand the impact of interloping gas, the correlation between window and density field (see e.g. [71, 72] for related investigations), and the precision of available cluster redshifts, though we expect these to be relatively minor issues.. In its nominal configuration, SO will detect 16,000 tSZ clusters, promising to get near the $M_{\rm{halo}}\sim 10^{14}\,M_{\odot}$ mark [73, 74]. This will improve further with CMB-S4, which is expected to detect between 45,000 and 140,000 clusters, depending on the final design choice [35]. While CMB measurements are the most promising way to detect clusters at high redshift, X-ray measurements from eRosita [75] promise to deliver a highly complete catalog of clusters at $z<0.5$ [35], with masses well below $M_{\rm{halo}}\sim 10^{14}\,M_{\odot}$.

### v3

@@ Abstract
The radial positions of galaxies inferred from their measured redshift appear distorted due to their peculiar velocities. We argue that the contribution from stochastic velocities — which gives rise to ‘Fingers-of-God’ (FoG) anisotropy in the inferred maps — does not lend itself to perturbative modelling already on scales targeted by current experiments. To get around this limitation, we propose to remove FoG using data-driven indicators of their abundance that are local in nature and thus avoid selection biases. In particular, we show that the scale where the measured power spectrum quadrupole changes sign is tightly anti-correlated with both the satellite fraction and the velocity dispersion, and can thus be used to select galaxy samples with fewer FoG. In addition, we show that excluding galaxies in haloes more massive than a given mass threshold can help to discard many of the most problematic galaxies. Such selection could be achieved in practice using maps of the thermal Sunyaev-Zel’dovich distortion of the cosmic microwave background frequency spectrum. These techniques could potentially improve reconstructions of the large-scale velocity and displacement fields from the redshift-space positions of galaxies. They may also extend the reach of perturbative models for galaxy clustering, though in practice we find only marginal gains when fitting one-loop EFTofLSS models to simulations with mitigated FoG due to the relevance of other effects entering at two-loop order.
@@ 1 Introduction
We also explore the possibility of removing galaxies living in massive halos, where their thermal velocities are large. This could be done using ancillary thermal Sunyaev-Zel’dovich [33] or X-ray data, which has the advatange of having a clean and largely local selection function. We showcase the potential of this approach and emphasize the powerful complementarity between upcoming galaxy redshift surveys and CMB experiments such as the Simons Observatory (SO) [34] and CMB-S4 [35].
@@ 4.1 Breakdown of contributions to $P_{\ell}(k)$
According to our discussion around figure 1, the contribution from central-satellite and satellite-satellite pairs potentially becomes non-perturbative when the quadrupole of their power spectrum goes negative. In order to determine exactly what scale this corresponds to in either case, and also to understand the relative importance of different contributions to the measured multipoles, we will use the Abacus mocks again, populated with galaxies using a fixed HOD model — concretely, the ‘baseline’ model of Ref. [46] with the best-fit value from analyzing BOSS CMASS data; the only changes we introduce is to set the completeness to $f_{\mathrm{ic}}=1$ and to set the velocity bias to zero to avoid additional modeling difficulties [^7] All in all, the HOD parameters are the following: $\log_{10}(M_{1}/h^{-1}\,M_{\odot})=14.10$, $\kappa=0.2$, $\alpha=1.12$, $\log_{10}(M_{\rm cut}/h^{-1}\,M_{\odot})=12.86$, $\log_{10}\sigma=-2.8$, $\alpha_{\rm c}=0$, $\alpha_{\rm s}=1$ and $f_{\rm ic}=1$.. This time, however, we will use the weights in equation (3.5) to disentangle contributions from central-central, central-satellite and satellite-satellite pairs.
@@ 4.2.2 An analysis of Abacus mocks with varying satellite fractions
Since the satellite fraction correlates strongly with FoG strength, these results illustrate the difficulty of accounting for FoG in perturbative models such as the EFTofLSS. However, the reader will have noted that even in the case where we have no satellites (and no peculiar velocity of the centrals relative to the halo) , cosmological constraints become biased by $k_{\rm max}=0.25\,h\,\rm{Mpc}^{-1}$. This is in qualitative agreement with earlier work in Ref. [15] which showed, using $N$-body simulations populated with emission-line galaxies following an HOD, that lowering the satellite fraction leads to only slightly more accurate constraints on cosmological parameters with no significant gains in terms of the $k$-reach of the one-loop theory. This could well be because higher-loop contributions independent of FoG are becoming relevant. Unlike FoG, these corrections may well be perturbative; for example, Ref. [64] showed recently that two-loop corrections to the redshift-space matter power spectrum at $z=0.5$ are significant already at $k=0.15\,h\,\rm{Mpc}^{-1}$. In the next section, we set out to disentangle the impact of FoG from these other, higher-loop effects.
@@ 4.2.3 An analysis of a manifestly perturbative data vector acted on by FoG
We now use the central value of these best-fit parameters to draw new central-central, central-satellite and satellite-satellite spectra from the EFT model and scale them by FoG damping factors as required; finally we linearly combine the terms weighting each one of them by the appropriate factors of $f_{\rm sat}$. The resulting power spectrum multipoles for the full sample — extracted from the redshift-space power spectrum of equation (4.7) using (2.5) — are shown in figure 8. We also compare them to measurements on the Abacus mocks, seeing that our fitting ensures good qualitative agreement. Our synthetic data vector thus contains a mixture of central and satellite contributions that is representative of the truth (insofar as the simulations are realistic), but crucially has no two-loop contributions besides those introduced by FoG.
@@
The more relevant question is therefore whether cosmological constraints are biased when FoG are present. This subsection addresses this point in the limit that our synthetic data vector is made up exclusively of satellites (we will consider the ‘full sample’ in the next subsection). To do this, we will once again sample the posterior probability on the parameters given the synthetic data, using the priors in table 2, with only slight modifications to the likelihood described in §4.2.2. We once again use the Effort emulator for the velocileptors model, including AP distortions. Although the data vector is known with no uncertainty, it is still important to assign the correct relative weight between different $k$’s when fitting. We thus use an analytic disconnected covariance described in appendix B, assuming there is no shot noise in our measurements. The high accuracy of the Effort emulator and the fact that our data vector is noiseless allow us to rescale the covariance to an effective volume of $1000\,\rm{Gpc}^{3}$ while keeping emulator error limited to at most $0.5\,\sigma$ at each $k$. Working with such a large effective volume lets us avoid projection effects and dramatically reduces the uncertainty in our determination of the inference biases associated with FoG. Also, with such a large volume, the prior becomes relatively unimportant [67], and changes to it do not significantly impact our results.
@@
It is instructive to investigate the direction in which FoG drive the EFT parameters. In figure 10, we show the marginalized constraints on the nuisance parameters corresponding to the inferences above. In addition to the posterior mean and 68% credible intervals, we overlay also the prior. The reader may recall that the prior on the counterterm parameters was chosen so as to ensure that the EFT contributions remain a small fraction of the linear-theory contribution at $k_{\rm max}$; more specifically, one prior standard deviation away from zero, $\alpha_{n}$ constitutes a 50% correction to the linear-theory prediction for the $n$-th moment [22]. From figure 10, we learn that when FoG are present, the preferred counterterm parameter values lie in regions that are disfavored by the prior. $\alpha_{2}$, in particular, lies more than one standard deviation away from the prior center already at $k_{\rm max}=0.1\,h\rm{Mpc}^{-1}$ whether or not the hexadecapole is included. This points to a breakdown of perturbativity roughly at the scale where the quadrupole crosses zero. Note, however, that all counterterms respond: this is because stochastic velocities introduce a number of counterterms into the theory, all but one of which are fully degenerate with the redshift-space counterterms.
@@
At this point, it is worth noting that the lack of two-loop effects in this toy model makes a direct comparison between figures 11 and 7 unwarranted. Nevertherless, we do note that the trend seen in figure 7 where going from $f_{\rm sat}=0$ to $f_{\rm sat}=0.11$ shifts $\Delta\omega_{c}$ and $\Delta H_{0}$ in the negative direction and $\Delta A_{s}$ in the positive is reproduced also in figure 11 as the satellite fraction is increased.
Once again, it is worth visualizing the constraints on nuisance parameters corresponding to these inferences. These are shown in figure 12. The striking growth in $b_{1}$ with the satellite fraction is due to the mock sample having a higher proportion of satellites, which have higher bias. More interestingly, the $\alpha_{n}$ counterterms absorb FoG contributions and become large compared to the standard deviation of the prior imposed on them, which suggests that the counterterms are no longer a small correction to the linear theory contribution. This is important for two reasons: first, it tells us that the perturbative expansion is breaking down; and second, in current data analyses, the priors imposed on the counterterms are typically chosen to be informative in order to avoid prior volume effects. For a given prior on $\alpha_{n}$, mischaracterizing the impact of FoG contributions to this parameter can effectively shrink the prior on contributions from other processes and potentially bias cosmological constraints.
@@ 5.1 Discarding the reddest galaxies
The connection we showed between the zero-crossing of the quadrupole and the prevalence of FoG suggests using this observable as a data-driven and model-independent way to select samples less affected by FoG. One could imagine defining a selection in galaxy color space, measuring the quadrupole, perturbing the selection, referring to the quadrupole again, and iterating this procedure until the tradeoff between minimizing the quadrupole zero-crossing and respecting some other constraints (e.g. the number density of galaxies) is optimized. Though investigating this approach in detail is beyond the scope of this particular paper, we will now explore what can be considered the first step in this iteration: discarding the reddest galaxies in a given sample. It is well known that red galaxies display stronger FoG [29, 30, 31, 32]. This is related to the fact that stochastic velocities originate primarily from galaxies living in the most massive halos, and such high density regions are predominantly populated by redder galaxies, whereas bluer ones tend to avoid such environments [68, 69]. Indeed, Ref. [70] (hereafter SYS21), showed that the zero-crossing of the quadrupole is very sensitive to the color of galaxies (see their figure 15), and, to a lesser extent, also to their luminosity (their figure 14). Concretely, they showed that the power spectrum quadrupole of the reddest 20% of BOSS LOWZ galaxies crosses zero at $k\approx 0.2\,h\,\rm{Mpc}^{-1}$, while the full sample has a crossing at $k\approx 0.3\,h\,\rm{Mpc}^{-1}$. This suggests that splitting samples in color and simply dropping the reddest quintile could help mitigate FoG. Since the color selection is ‘local’ (in the sense that it is sensitive to the properties of individual galaxies rather than the large-scale structure of the Universe), this selection should not introduce the kinds of modeling complications that plague alternative approaches to FoG removal.
When trying to study the impact of such color-based selection, we face the difficulty that the Abacus simulations do not mock up the color of galaxies. In spite of this, we will still be able to approximate different color subsamples somewhat roughly by ensuring that they reproduce the power spectra of the color splits reported in SYS21. Let us now explain how this can be done.
@@
Figure 13: Power spectrum multipoles approximating the effect of removing the reddest 20% of galaxies from a BOSS-LOWZ-like sample. (Unlike in other plots in this work, we do not subtract shot noise from the monopole.) By reweighting galaxies to have an effective $f_{\rm sat}=0.3$ ($f_{\rm sat}=0.15$), we mock up a sample with a zero-crossing of the quadrupole at $k=0.21$ ($k=0.31$) resembling that of the reddest quintile (full sample) of BOSS LOWZ [70]. From this we can predict the spectra of the remaining 80% of galaxies after the reddest quintile is discarded, which has a zero-crossing at a higher $k=0.36$. This suggests that the remaining sample has a significantly smaller scale of FoG non-linearity.
@@
Figure 14: Marginalized posterior mean and 68% credible intervals on cosmological parameters as a function of $k_{\rm max}$ before (purple) and after (yellow) reweighting galaxies in a way that mimics discarding the reddest quintile of mock galaxies resembling the BOSS LOWZ selection.
@@ 5.2 Masking galaxy clusters and groups
In §3, we showed that FoG are sourced primarily by galaxies living in the most massive halos, where deep gravitational potential wells source large virial velocities. Before quantifying the extent to which removing these galaxies can improve cosmological constraints from the full-shape of the galaxy power spectrum, let us briefly comment on what ancillary data would enable such a selection.
@@
Massive galaxy clusters produce strong X-ray emission and Sunyaev-Zel’dovich (tSZ) distortions of the cosmic microwave background frequency spectrum. tSZ and X-ray intensity are tightly correlated with halo mass, so masking on these observables — for example, in regions with SNR of 5 or higher — could potentially remove the galaxies with the strongest peculiar velocities [^14] There is likely a connection between what we propose to do and density-split statistics or marked power spectra, though we claim that our selection is more local in nature as it is guided by halo mass. Our approach will thus produce populations with clustering properties that are easier to model. The modeling will also be made easier by the fact that we typically identify — and discard — a very small number of objects.. The selection and removal of these galaxies would thus be very clean and depend only on halo mass and crucially not on galaxy overdensity or velocity [^15] This is due in large part to the fact that structures that can be detected on an individual basis in current or upcoming Compton $y$-maps are all in the one-halo regime, with two-halo contributions to the tSZ effect only becoming evident after stacking over very many objects (the impact of two-halo contributions to the $y$-maps could be checked on a case-by-case basis using hydrodynamical simulations). Moreover, these objects are typically identified only after high-pass filtering the Compton $y$-maps to eliminate any long-wavelength modes tracing the cosmic web. This means that the key discerning property in this selection is halo mass. The resulting galaxy sample will simply have a different (smaller) galaxy bias than the initial population; bias can of course be marginalized over in the standard ways, at least in the limit applicable here where only a relatively small fraction of the galaxies are masked.. (Below, we will in fact verify that thresholding on halo mass does not introduce any noticeable biases on parameter inference; much on the contrary, it helps recover more accurate cosmological parameters.) As an example, suppose we are to analyze a galaxy redshift survey that fully overlaps with a complementary CMB survey. The tSZ signal being independent of redshift, we could in principle have a complete census of all the clusters above a certain mass threshold, and discard galaxies living within them [^16] Future work will be required to understand the impact of interloping gas, the correlation between window and density field (see e.g. [71, 72] for related investigations), and the precision of available cluster redshifts, though we expect these to be relatively minor issues.. In its nominal configuration, SO will detect 16,000 tSZ clusters, promising to get near the $M_{\rm{halo}}\sim 10^{14}\,M_{\odot}$ mark [73, 74]. This will improve further with CMB-S4, which is expected to detect between 45,000 and 140,000 clusters, depending on the final design choice [35]. While CMB measurements are the most promising way to detect clusters at high redshift, X-ray measurements from eRosita [75] promise to deliver a highly complete catalog of clusters at $z<0.5$ [35], with masses well below $M_{\rm{halo}}\sim 10^{14}\,M_{\odot}$.